OpenStudy (anonymous):

What is the simplified form of each expression? -√196 A. 98 B. –14 C. 14 D. 392

OpenStudy (anonymous):

B

OpenStudy (anonymous):

- 14 X -14 = 196

OpenStudy (venomblast):

it B. the square root of 196 is 14. since you have the negative out side the radical, it becomes negative. (remeber the negative outside the radical is actually a -1) after squaring rooting 196, u get 14. so -1 x 14 = -14

OpenStudy (anonymous):

YUP can anyone help me on my problem please Its a picture

OpenStudy (anonymous):

i dont under Algebra I Part

OpenStudy (anonymous):

sure i can true.

OpenStudy (anonymous):

Oh, I can help with that one. Ready to work together on it?

OpenStudy (anonymous):

Yeah sure

OpenStudy (anonymous):

The easiest way to solve a multiple choice problem like that is to plug in the x values and try to find the answer through elimination.

OpenStudy (anonymous):

OKay

OpenStudy (anonymous):

I can tell you that plugging in 0 won't help because they all turn into -3. So, let's do 1 for each one. A. y = –x – 3 B. y = x – 3 C. y = 2x − 3 D. y = ­­–2x + 3 Just to make sure you understand, where do we plug in the "1" that we're going to plug in, and what value do we want the answer to be?

OpenStudy (anonymous):

i dont understand Algebra I Part and i have home work.

OpenStudy (anonymous):

* (Well I lied when I said they all turn to -3 when you plug in 0; D turns into +3, so we know D is wrong already! Yay!)

OpenStudy (anonymous):

YAY

OpenStudy (anonymous):

Alright, so where do you think we plug in the 1 when we are testing equation A? A. y = –x – 3

OpenStudy (anonymous):

Where the x is?

OpenStudy (anonymous):

x-actly! (Yay, puns!) And what happens when we do: y = -(x) - 3 (make sure you never forget that negative in front of the x!)

OpenStudy (venomblast):

answer is C

OpenStudy (venomblast):

you plug in for x for each and try to get y.

OpenStudy (anonymous):

AWWE your so smart thank you

OpenStudy (venomblast):

to who you saying thank you?

OpenStudy (anonymous):

You have the same potential to answer the questions; you've just gotta try to solve 'em. lol

OpenStudy (anonymous):

you both are so smart at math .

OpenStudy (anonymous):

haha thanks

OpenStudy (anonymous):

But Gabriella, just to make sure: do you understand HOW to solve a similar problem if one appeared on a test?

OpenStudy (anonymous):

Yes :)

OpenStudy (anonymous):

Ah alright, that's great then!